# Comprehensive Cognitive LLM Agent for Smartphone GUI Automation

### v2

@@ 5.3 Future Actions
This section considers a more challenging setting of next actions prediction. The task is much more non-trivial as the agent only receives the environment state , without the perception of future states . Thus, predicting future actions involves harder reasoning, planning, and environment simulation. Table 8 shows the results with . Although the next action can be predicted with 70.96% accuracy, predicting the following actions without environmental feedback remains to be improved.
@@ 6 Conclusion
This paper proposes CoCo-Agent, an LLM-based autonomous agent for smartphone GUI. Our method facilitates comprehensive cognition with exhaustive perception and reliable action response. CoCo-Agent is enhanced with two approaches. Comprehensive environment perception (CEP) integrates GUI perception, and conditional action prediction (CAP) enables a decomposition of complex commands. CoCo-Agent achieves SOTA performance in experiments on two GUI datasets. Further analysis shows that the agent learns the behavior patterns in smartphone GUI. The significance of the proposed enhancements is also verified. We discuss the unbalanced category distributions and the underestimation of agent performance. Our work reveals the promising capabilities of LLMs as autonomous agents, especially for complex environment perception and action response.
@@
We acknowledge the limitations of this work. (i) Resource consumption. The data is of a large amount. The training process costs computational resources compared to the zero-shot methods. Whereas, the training process only needs to be conducted once. And our unified model achieves the SOTA. Our future work will seek a balance between the training cost and performance. (ii) More difficult settings. As is shown in Section 5.3, the predictions of future action remain to be improved. The ultimate goal is to empower the agent to operate the full episodes in a simulated GUI environment or a realistic device.
@@ Appendix A Implementation Details
Experiment Details. For AITW, we follow the training and test set splits on General, Install, Google Apps, and Web shopping of Rawles et al. (2023), and use the splits on Single of Zhang and Zhang (2023), which is more available. Each subset is split for training, validation, and test sets by 8:1:1. The first 1000 samples of the test set are used as the validation sets. The reported scores are the results of the full test sets. The maximum length of action history is 8-lastest action. The model is trained for {8,10,12} epochs, with a learning rate of 2e-5. The batch size is set to {12,16} per device. All experiments are conducted on 4 Nvidia A800 GPUs.
@@ Appendix B Examples
We show examples of a full episode in Figure 2. Examples of the randomness in AITW are illustrated here. Figure 3 is a case where the predicted action leads to the same situation but does not match the label (i, in Section 5.4.2). Figure 4 is a case where the predicted action is different from the label but is still reasonable for the goal (ii, in Section 5.4.2). The target items on the screen are highlighted by blue boxes for label actions and orange boxes for predicted actions.

### v3

@@ 5.3 Future Actions
This section considers a more challenging setting of next actions prediction. The task is much less trivial as the agent only receives the environment state , without the perception of future states . Thus, predicting future actions involves harder reasoning, planning, and environment simulation. Table 8 shows the results with . Although the next action can be predicted with 70.96% accuracy, predicting the following actions without environmental feedback remains to be improved.
@@ 6 Conclusion
This paper proposes CoCo-Agent, an MLLM-based autonomous agent for smartphone GUI. Our method facilitates comprehensive cognition with exhaustive perception and reliable action response. CoCo-Agent is enhanced with two approaches. Comprehensive environment perception (CEP) integrates GUI perception, and conditional action prediction (CAP) enables a decomposition of complex commands. CoCo-Agent achieves SOTA performance in experiments on two GUI datasets. Further analysis shows that the agent learns the behavior patterns in smartphone GUI. The significance of the proposed enhancements is also verified. We discuss the unbalanced category distributions and the underestimation of agent performance. Our work reveals the promising capabilities of MLLMs as autonomous agents, especially for complex environment perception and action response.
@@
We acknowledge the limitations of this work. (i) Resource consumption. The data is of a large amount. The training process costs computational resources compared to the zero-shot methods. Whereas, the training process only needs to be conducted once. And our unified model achieves the SOTA performance. Efficient LLM training or inference methods can improve the balance between resource cost and performance. (ii) More complex settings. As is shown in Section 5.3, the predictions of future action remain to be improved. The ultimate goal is to empower the agent to operate the full episodes in a simulated GUI environment or a realistic device.
Future Work
The following directions for future work are proposed based on our study. First, GUI agents require generalization ability to support new instructions, new applications, or even new operating systems. The effectiveness of CoCo-Agent in task adaptation across domains of AITW has been preliminarily validated. Second, GUI agents require improved multimodal training strategies. Multimodal LLMs can be strengthened by integrating GUI perception into vision-language pre-training or instruction tuning. Third, GUI agents require comprehensive measurements. The measurements can be improved to reflect different paths for the same goal in practical scenarios. Last but not least, there is still room for performance improvement (79.05% for AITW and 88.27% for META-GUI) and future action reasoning and planning for full-episode prediction.
Ethics Statement
This section presents ethics statements in the following aspects. (i) Data Privacy. The research dataset AITW Rawles et al. (2023) has declared control over Personal Identifiable Information. The instructions contain no toxic, biased, or misleading content. The META-GUI is based on a task-oriented dialogue dataset, SMCalFlow Andreas et al. (2020), crowdsourced on Amazon Mechanical Turk with qualification requirements. CoCo-Agent does not rely on LLM APIs, preserving privacy data from leakage to LLM companies. (ii) System security. Compared to the command-line interface (CLI), the GUI is more interpretable and controllable. Since GUI actions follow human behavior, security considerations can better align with human-oriented mechanisms, which are already implemented in existing GUIs for operating systems. However, the potential risks of GUI agents have yet to be well explored Yuan et al. (2024); Yang et al. (2024). (iii) Potential social impacts. On the one hand, GUI automation can facilitate efficiency and save human resources for more high-level work. On the other hand, malicious actors could abuse GUI agents to achieve undesirable purposes. For practical applications of GUI agents, platforms may need to update detection, authorization, and governance mechanisms to control potential risks for social impacts.
@@ Appendix A Implementation Details
Experiment Details. For AITW, we follow the training and test set splits on General, Install, Google Apps, and Web shopping of Rawles et al. (2023), and use the splits on Single of Zhang and Zhang (2023), which is more available. Each subset is split for training, validation, and test sets by 8:1:1. The first 1000 samples of the test set are used as the validation sets. The reported scores are the results of the full test sets. For META-GUI, we follow the original dataset splits. The maximum length of action history is 8-lastest action. The model is trained for {8,10,12} epochs, with a learning rate of 2e-5. The batch size is set to {12,16} per device. All experiments are conducted on 4 Nvidia A800 GPUs.
@@ Appendix B Examples
We show examples of a full episode in Figure 2. Examples of the randomness in AITW are illustrated here. Figure 3 is a case where the predicted action leads to the same situation but does not match the label (i, in Section 5.4.2). Figure 4 is a case where the predicted action is different from the label but is still reasonable for the goal (ii, in Section 5.4.2). Blue boxes highlight the target items on the screen for label actions and orange boxes for predicted actions.
Appendix C Future Work
The following directions for future work are proposed based on our study. First, GUI agents require generalization ability to support new instructions, new applications, or even new operating systems. The effectiveness of CoCo-Agent in task adaptation across domains of AITW has been preliminarily validated. Second, GUI agents require improved multimodal training strategies. Multimodal LLMs can be strengthened by integrating GUI perception into vision-language pre-training or instruction tuning. Third, GUI agents require comprehensive measurements. The measurements can be improved to reflect different paths for the same goal in practical scenarios.